# DP-BREM: Differentially-Private and Byzantine-Robust Federated Learning with Client Momentum

### v2

@@ Appendix A Proof of Lemma 2 (Aggregation Sensitivity)
Appendix B Proof of Theorem 1 (Privacy Analysis)
Proof.
Also, due to the client-level sampling (i.e., each client was selected by the server w.p. $q$) and record-level sampling (i.e., each record was selected by client $\mathsf{C}_{i}$ w.p. $p_{i}$), the overall sampling rate is $qp_{i}$. Then, by applying the privacy accountant of Gaussian DP (GDP) [11] shown in Lemma 10 (see more details in Appendix F), DP-BREM satisfies $\mu_{i}$-GDP with $\mu_{i}$ shown in (16). Finally, by converting $\mu_{i}$-GDP to $(\epsilon_{i},\delta)$-DP via Lemma 9, we get (15), which finishes the proof. ∎
Appendix C Proof of Theorem 2 (Aggregation Error)
Proof.
Our proof heavily relies on several useful lemmas shown in Appendix E, where Lemma 4 splits the L2-norm of summation of vectors into weighted summation of vectors’L2-norm, and Lemma 5 provides the optimal strategy to choose these weights.
where
=$\overset{(\mathsf{b})}{=}[\rho R/(R+c)+\phi]^{2}$
$\leqslant|\mathcal{H}_{t}|\cdot[\rho R/(R+c)+\phi]^{2}$
where $(\mathsf{a})$ used the assumption that $\{\bar{\bm{m}}_{t,i}\}_{i\in\mathcal{H}_{t}}$ are independent, then the random variables $\{\bm{y}_{i}\}_{i\in\mathcal{H}_{t}}$ are also independent; $(\mathsf{b})$ used contractivity of a clipping (projection) step. Therefore,
Putting into Together. Combining all terms, we have
$\qquad\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{\mu}\|^{2}$
The Final Result. On the other hand, we have
$\qquad\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{m}_{t}^{*}\|^{2}$
where $(\mathsf{a})$ is obtained by using Lemma 4 for any $\gamma\in(0,1)$; $(\mathsf{b})$ is obtained from (C) and (C), where $\Phi$ is defined in (C); $(\mathsf{c})$ is obtained by taking $\gamma=\frac{\Phi}{\Phi+\frac{\rho}{\sqrt{|\mathcal{H}|}}}$. Furthermore, if we assume $\phi\leqslant O(\rho)$ and $\tau\leqslant O(\rho)$, we can rewrite (C) as the following version
which finishes the proof of Theorem 2. ∎
Appendix D Proof of Theorem 3 (Convergence Rate)
Proof.
Comparing with the aggregation error of $O(\rho^{2}|\mathcal{B}|/n)$ (ignoring constants and higher order terms) in [22, Lemma 9], our aggregation error shown in (19) replaces the term $|\mathcal{B}|$ by $|\mathcal{B}|+\sqrt{d}\sigma/q$, which means a slower convergence due to DP noise. Then, following the result in [22, Theorem VI] and its informal version in (5), we get the convergence rate of our algorithm as in (20). Note that our aggregation utilizes a client-level sampling rate $q$, i.e., approximate $nq$ clients participate in the aggregation for one iteration. We need to replace the term of $\frac{1}{n}$ in (5) by $\frac{1}{nq}$ in (20). ∎
Appendix E Useful Lemmas
Lemma 4.
For any positive real values $\alpha_{1},\cdots,\alpha_{K}\in\mathbb{R}^{+}$ and any $d$-dimensional vectors $\bm{x}_{1},\cdots,\bm{x}_{K}\in\mathbb{R}^{d}$, the following inequality holds
$\left\|\sum\nolimits_{k=1}^{K}\bm{x}_{k}\right\|^{2}\leqslant\left(\sum\nolimits_{k=1}^{K}\alpha_{k}\right)\cdot\left(\sum\nolimits_{k=1}^{K}\frac{\|\bm{x}_{k}\|^{2}}{\alpha_{k}}\right)$
[…]
Proof.
Denote $x_{ki}$ as the $i$-th element of the vector $\bm{x}_{k}$, then we have
where the inequality is caused by Cauchy-Schwarz inequality. ∎
Lemma 5.
Consider the following optimization problem
$f^{*}=\min_{x_{1},\cdots,x_{K}}~{}~{}\sum\nolimits_{k=1}^{K}\frac{c_{k}}{x_{k}},\qquad\text{such that}\quad x_{k}>0,\quad\sum\nolimits_{k=1}^{K}x_{k}=1$
Proof.
Appendix F Gaussian Differential Privacy (GDP)
@@
In this paper, we use $\mu$-GDP as our primary privacy accountant method due to its good property on composition and accountant of privacy amplification in Lemma 10, and then convert the result to $(\epsilon,\delta)$-DP via Lemma 9. We note that other privacy accountant methods, such as moments accountant [1] and Rényi DP (RDP) [33], are also applicable to the proposed scheme and theoretical analysis, but might lead to suboptimal results.
Appendix G Preliminaries for Crypto Primitives
@@
Appendix H Detailed Steps of DP-BREM+ in Figure 2
[…]
Appendix I Proof of Theorem 4 (Security Analysis)
@@
Appendix J Supplements of Experiments
Figure 8: MNIST: Iteration curve with privacy budget $\epsilon=3$ and Byzantine clients $\delta_{B}=30\%$ over $n=100$ clients.
Figure 9: CIFAR-10: Iteration Curve with privacy budget $\epsilon=4$ and Byzantine clients $\delta_{B}=15\%$ over $n=100$ clients.
J.1 Experimental Setup
Datasets (non-IID) and Model Architecture. We use two datasets for our experiments: MNIST [25] and CIFAR-10 [24], where the default value of the number of total clients is $n=100$. For MNIST dataset, we use the CNN model from PyTorch example [^4] https://github.com/pytorch/opacus. For CIFAR-10 dataset, we use the CNN model from the TensorFlow tutorial [^5] https://www.tensorflow.org/tutorials/images/cnn, like the previous works [49, 30]. To simulate the heterogeneous data distributions, we make non-i.i.d. partitions of the datasets, which is a similar setup as [49] and is described below:
3) LF (label-flipping). The attacker modifies the labels of all examples of corrupted clients’ data and trains a new model with multiple iterations, then uses model replacement strategy [2] to enhance the impact on the global model.
5) DP-LFH. The baseline (shown in Section 3.2) directly combines DP-SGD based momentum with LFH. Each client adds DP noise to the local gradient, and then computes the local momentum that will be aggregated with centered clipping by the server.
J.2 Parameters Setting
Basic Parameters.
  • Total number of iterations $T$: 1000 for MNIST; 2000 for CIFAR-10
  • Learning rate $\eta_{t}$: For MNIST datasets, $\eta_{t}$ is linearly reduced from 0.1 to 0.01 w.r.t. iterations. For CIFAR-10 dataset, $\eta_{t}$ is linearly reduced from 0.05 to 0.0025 w.r.t. iterations.
DP-related Parameters.
  • Record-level sampling rate $p_{i}$: 0.05 for all $i$
  • Record-level clipping bound $R$: linearly reduced from $R_{0}$ to $0.3R_{0}$ w.r.t. iterations. Note that in Figure 7, the different value of $R$ in x-axis is the value of the above $R_{0}$. For MNIST, we set $R_{0}=10$ by default, but $R_{0}=5$ only for the case of $\epsilon=1$ in Figure 5. For CIFAR-10, we set $R_{0}=20$ by default, but $R_{0}=15$ only for the case of $\epsilon=2$ in Figure 6.
  • Privacy parameter $\delta$ in DP: $10^{-6}$
  • Noise multiplier $\sigma$: For MNIST (with $T=1000$ and each client has $|\mathcal{D}|_{i}=60000/100=600$ examples), $\sigma\in\{0.15,0.06,0.029,0\}$ for $\epsilon\in\{1,3,8,\text{inf}\}$. For CIFAR-10 (with $T=2000$ and each client has $|\mathcal{D}|_{i}=50000/100=500$ examples), $\sigma\in\{0.14,0.077,0.042,0\}$ for $\epsilon\in\{2,4,9,\text{inf}\}$.
Robustness-related Parameters.
  • Client-level clipping bound $C$ (only for DP-BREM and DP-LFH): linearly reduced from $C_{0}$ to $0.3C_{0}$ w.r.t. iterations, where $C_{0}=1$ for MNIST, and $C_{0}=5$ for CIFAR-10.
  • Momentum parameter $\beta=0.9$
J.3 More Experimental Results
Iteration Curve. Figures 8 and 9 show how the accuracy changes with the training iterations in MNIST (with total iteration $T=1000$) and CIFAR-10 (with $T=2000$), respectively. Due to the existence of Byzantine attacks, the iteration curve is not as smooth as in the attack-free case.
Appendix K Other Related Work
@@
Private and Byzantine-Robust FL. Recently, some works tried to simultaneously achieve both privacy and robustness of FL. He et al. [20] proposed a Byzantine-resilient and privacy-preserving solution, which makes distance-based robust aggregation rules (such as Krum [5]) compatible with secure aggregation via MPC and secret sharing. So et al. [40] developed a similar scheme based on Krum, but rely on different cryptographic techniques, such as verifiable Shamir’s secret sharing and Reed-Solomon code. Velicheti et al. [42] achieved both privacy and Byzantine robustness via incorporating secure averaging among randomly clustered clients before filtering malicious updates through robust aggregation. However, these works only ensure the security of the aggregation step and do not achieve DP for the aggregated model.

### v3

@@ Appendix H Proof of Lemma 2 (Aggregation Sensitivity)
Appendix I Gaussian Differential Privacy (GDP)
@@
In this paper, we use $\mu$-GDP as our primary privacy accountant method due to its good property on composition and accountant of privacy amplification and then convert the $\mu$-GDP result to $(\epsilon,\delta)$-DP. We note that other privacy accountant methods, such as moments accountant [1] and Rényi DP (RDP) [34], are also applicable to the proposed scheme and theoretical analysis, but might lead to suboptimal results.
Appendix J Preliminaries for Crypto Primitives
@@
Figure 6: Iteration Curve for three datasets.
Appendix K Proof of Theorem 4 (Security Analysis)
@@
Appendix L More Experimental Results
Iteration Curve. Figure 6 shows how the accuracy changes with the training iterations across three datasets. The CIFAR-10 dataset, being more complex, has a larger total training iterations ($T=2000$), while the other two datasets have $T=1000$. The presence of Byzantine attacks results in less smooth iteration curves compared to the attack-free case.
Appendix M Other Related Work
@@
Private (non-DP) and Byzantine-Robust FL. Recently, some works tried to simultaneously achieve both privacy and robustness of FL. He et al. [21] proposed a Byzantine-resilient and privacy-preserving solution, which makes distance-based robust aggregation rules (such as Krum [5]) compatible with secure aggregation via MPC and secret sharing. So et al. [41] developed a similar scheme based on Krum, but rely on different cryptographic techniques, such as verifiable Shamir’s secret sharing and Reed-Solomon code. Velicheti et al. [43] achieved both privacy and Byzantine robustness via incorporating secure averaging among randomly clustered clients before filtering malicious updates through robust aggregation. However, these works only ensure the security of the aggregation step and do not achieve DP for the aggregated model.